# Dynamic Routing Between Experts: A Data-Efficient Approach to Continual Learning in Vision-Language Models

### v2

@@ Appendix A Dataset description and sizes
COCO-Caption dataset consists of over 330,000 images, each annotated with five human-generated captions, originally designed for image captioning tasks. We use an adapted split, which includes 566,747 image-text pairs for training and 5,000 image-text pairs for testing. This provides a rich multimodal resource for exploring the relationships between visual content and natural language descriptions. The performance is reported using the CIDEr (Vedantam, Lawrence Zitnick, and Parikh 2015) metric, which evaluates the quality of generated captions based on their similarity to reference captions.
SNLI-VE is derived from the Stanford Natural Language Inference (SNLI) (Bowman et al. 2015) corpus, SNLI-VE (Visual Entailment) extends this dataset by including visual contexts. It consists of 529,527 image-text pairs for training, with a test set of 2,000 pairs. These pairs are categorized into three labels: (a) entailment, (b) neutral, and (c) contradiction, allowing models to evaluate their ability to reason about the relationships between visual and textual information. We use accuracy as the evaluation metric for this task.
@@
Benchmarks
In addition to evaluating metrics on fine-tuning tasks, we conduct an in-depth analysis of the fine-tuned model using well-established benchmarks to assess the extent of forgetting introduced during model training. For this analysis, we utilize the MMBench (Liu et al. 2025), ChartQA (Masry et al. 2022b), and DocVQA (Mathew, Karatzas, and Jawahar 2021a) benchmarks, each targeting distinct evaluation objectives.
@@
Appendix B Related work
Continual Learning
Experience replay methods address CF by maintaining a small buffer of previously observed data, which is reused to retrain the model when learning new tasks. Although these methods theoretically pose a risk of looser generalization bounds, they are widely valued for their simplicity, stability, and robust performance, even with limited episodic memory buffers (Chaudhry et al. 2018; Rolnick et al. 2019). This practicality has made replay-based approaches a popular choice in continual learning scenarios (Rebuffi et al. 2017; Lopez-Paz and Ranzato 2017). The Instruction-based Continual Learning (InsCL) method (Wang et al. 2024c) extended this idea by introducing dynamic replay, which selects prior data based on task similarity, measured using the Wasserstein Distance with task-specific instructions. However, our experiments demonstrate that this approach results in forgetting earlier tasks, highlighting its limitations in retaining comprehensive task knowledge.
Regularization methods tackle CF by adding a penalty term to the loss function, which discourages significant changes to model parameters that are critical for previously learned tasks. The key principle is to balance retaining prior knowledge with acquiring new information, typically controlled by a regularization coefficient. Notable examples include Elastic Weight Consolidation (EWC) (Kirkpatrick et al. 2017) and Learning without Forgetting (LwF) (Li and Hoiem 2017). EWC constrains significant changes to parameters deemed crucial for earlier tasks, while LwF employs a knowledge distillation loss (Hinton et al. 2015) to preserve performance on previous tasks. Despite their effectiveness, these methods often increase memory consumption and can be challenging to train due to the difficulty in achieving an optimal balance between old and new tasks. A routing-based approach offers a promising alternative, mitigating these issues by providing greater flexibility in balancing tasks while optimizing memory efficiency.
Federated Learning
Federated Learning (FL) has become a pivotal paradigm in distributed machine learning, enabling multiple decentralized modules to collaboratively train a shared model while keeping raw data local. This framework addresses critical privacy and data sovereignty concerns by aggregating model parameters rather than exchanging sensitive data (McMahan et al. 2017). However, FL encounters several challenges, including handling non-IID (non-independent and identically distributed) data across modules, mitigating communication inefficiencies, and ensuring equitable performance across diverse modules (Kairouz et al. 2021; Li et al. 2020). Techniques like FedAvg (McMahan et al. 2017) and its adaptations have sought to address these issues by optimizing communication protocols and improving robustness to heterogeneous data distributions.
CL within the FL paradigm introduces further complexities due to the dynamic nature of non-stationary data distributions (Yoon, Yang, and Hwang 2021). CF becomes particularly pronounced as models need to adapt to new tasks while retaining prior knowledge without direct access to historical data. Federated continual learning (FCL) approaches tackle these challenges by leveraging methods such as knowledge distillation (Zhao et al. 2022), experience replay (Shoham, Rotem, and Ben-Ari 2022), and personalized model training (T Dinh, Tran, and Nguyen 2020). These strategies strive to balance global knowledge retention with local task adaptation, yet scalability and robustness remain open challenges. Modular machine learning offers a scalable solution to this interplay by enabling task-specific pathways, where only essential modular updates are shared rather than entire model parameters. This reduces communication overhead while preserving privacy. However, the primary challenge with routing lies in training these task-specific pathways effectively, which we identify as an important direction for future research.
Appendix C Routing patterns
@@
Appendix D Routing in cross-lingual setting
In this section, we investigate whether routing can enable cross-lingual transfer—specifically, whether a routing model can solve a task in one language by combining a LoRA module trained on that language with another LoRA module trained on a specialized task in English. To evaluate this, we focus on the MGSM dataset (Shi et al. 2022), a multilingual math-based reasoning benchmark. We construct two expert models: a Math Expert, trained on the OrcaMath dataset (Mitra et al. 2024), and a Chinese Language Expert, trained on the xP3mt dataset (Muennighoff et al. 2022).
@@ Appendix D Routing in cross-lingual setting
Appendix E Training details

### v3

@@ Appendix A Dataset description and sizes
COCO-Caption dataset consists of over 330,000 images, each annotated with five human-generated captions, originally designed for image captioning tasks. We use an adapted split, which includes 566,747 image-text pairs for training and 5,000 image-text pairs for testing. This provides a rich multimodal resource for exploring the relationships between visual content and natural language descriptions. The performance is reported using the CIDEr Vedantam et al. (2015) metric, which evaluates the quality of generated captions based on their similarity to reference captions.
SNLI-VE is derived from the Stanford Natural Language Inference (SNLI) Bowman et al. (2015) corpus, SNLI-VE (Visual Entailment) extends this dataset by including visual contexts. It consists of 529,527 image-text pairs for training, with a test set of 2,000 pairs. These pairs are categorized into three labels: (a) entailment, (b) neutral, and (c) contradiction, allowing models to evaluate their ability to reason about the relationships between visual and textual information. We use accuracy as the evaluation metric for this task.
@@
A.1 Benchmarks
In addition to evaluating metrics on fine-tuning tasks, we conduct an in-depth analysis of the fine-tuned model using well-established benchmarks to assess the extent of forgetting introduced during model training. For this analysis, we utilize the MMBench Liu et al. (2024b), ChartQA Masry et al. (2022), and DocVQA Mathew et al. (2021) benchmarks, each targeting distinct evaluation objectives.
@@ A.1 Benchmarks
Appendix B Routing patterns
@@
Appendix C Routing in cross-lingual setting
In this section, we investigate whether routing can enable cross-lingual transfer—specifically, whether a routing model can solve a task in one language by combining a LoRA module trained on that language with another LoRA module trained on a specialized task in English. To evaluate this, we focus on the MGSM dataset Shi et al. (2022), a multilingual math-based reasoning benchmark. We construct two expert models: a Math Expert, trained on the OrcaMath dataset Mitra et al. (2024), and a Chinese Language Expert, trained on the xP3mt dataset Muennighoff et al. (2023).
@@ Appendix C Routing in cross-lingual setting
Appendix D Training details